OpenStudy (anonymous):

What constant term should be used to complete the square? x 2 - 5x + _____ = 7 -

OpenStudy (wolf1728):

To complete the square, take the coefficient of x (which is -5) , divide it by 2 then square it, then add it to both sides of the equation. So, minus 5 becomes 6.25 and the equation becomes x 2 - 5x + 6.25 = 7 +6.25 If you need to solve the equation it's pretty well set up now.

OpenStudy (anonymous):

i got but it does not match any answers x=12(5+53−−√),12(5−53−−√) x≈6.14005494464,−1.14005494464

OpenStudy (anonymous):

is wolf still there please can u still help.i will give a medal

OpenStudy (wolf1728):

Hello wildfire

OpenStudy (anonymous):

hey did u see my answer it does not match anything that they say??

OpenStudy (wolf1728):

I'm back - I got "bounced" off the site - happens a LOT.

OpenStudy (anonymous):

whew so what do u think

OpenStudy (wolf1728):

Are you showing all the answers?

OpenStudy (anonymous):

-

OpenStudy (anonymous):

49/25, 25/4, 4/25, 49/25

OpenStudy (wolf1728):

Wow - this board is REALLY screwy right now. I guess we should solve the equation Square root of 13.25 =4.0311288741 and -4.0311288741 x = 4.0311288741 + 2.5 x = -4.0311288741 + 2,5

OpenStudy (wolf1728):

x = 6.5311288741 x = -1.5311288741

OpenStudy (anonymous):

so which ine would it be that is what i get but as u can see i have only those choices

OpenStudy (wolf1728):

If I can type an answer the solutions are 6.1400549446 -1.1400549446 I got the square root wrong. These look as though they are solutions yoy have.

OpenStudy (wolf1728):

I trust you can see that right?

OpenStudy (anonymous):

yes i have been stuck on this so long i quess atthis point i will guess if i get it incorrect i will learn from my mistake

OpenStudy (wolf1728):

Okay my last posting has the correct answers - 6.1400549446 -1.1400549446 Which are on your list

OpenStudy (wolf1728):

I imagine you aren't answering because this board is acting equally crazy with you as it is with me.

OpenStudy (wolf1728):

Well gonna get going - good luck